I have no idea why I got a certain answer. This is my problem: 5/6x+30=x then I multiplied both by 6 and added them together to get 185. Am I correct? If so, how does that answer work?
Answer by nyc_function(2741)

Nice try but you needed to multiply all terms by 6x.

If your question is (5/6x) + 30 = x, then your answer will be positive and negative for x.

(5/6x) + 30 = x

6x * (5/6x) + 30 * 6x = x * 6x

5 + 1800x = 6x^2

6x^2 - 1800x -5 = 0

This is quadratic equation.

Using the quadratic formula (I assume that you know how to use this formula),
I got the following two answers for x.

x = [90 + sqrt{8130}]/6 and x = [90 - sqrt{8130}]/6

NOTEL sqrt = square root for short

If you need to learn how to use the quadratic formula, watch this video clip.
http://www.youtube.com/watch?v=UgIF7Q7lAuc